# On stable assignments generated by choice functions of mixed type

### v1

@@ 4 Rotations
(i) for each $v\in J$, the value $\alpha(A_{v})-\beta(B_{v})$ is zero;
@@
Definition 4. Let us call an edge $e=fw$ active for $F$ if $f\in F^{+}$ and $e\in D_{f}$, and active for $W$ if $w\in W^{+}$ and $e\in H_{w}$. The subgraph of $G$ induced by the set of active edges is called the active graph and denoted as $\Gamma=(V_{\Gamma},E_{\Gamma})$. The directed counterpart of the active graph, denoted as $\overrightarrow{\Gamma}=(V_{\Gamma},\overrightarrow{E}_{\Gamma})$, is obtained by replacing each edge $e=fw\in E_{\Gamma}$ by the directed edge $\overrightarrow{e}=(f,w)$ if $e\in D_{f}$, and by the directed edge $\overrightarrow{e}=(w,f)$ if $e\in H_{w}$. A (weak) component of $\Gamma$ (or $\overrightarrow{\Gamma}$) is a maximal subgraph $K$ in it such that any two vertices of $K$ are connected by a (not necessarily directed) path.
@@
Proof  Take as $A$ the set $\cup(D_{f}\colon f\in F_{K})$ and take as $B$ the set $\cup(H_{w}\colon w\in W_{K})$, where $F_{K}:=F\cap V_{K}$ and $W_{K}:=W\cap V_{K}$. Property ((4.3)) implies that $A,B$ obey condition ((4.12)). Then, by Lemma 4.3, there exist $\alpha:A\to{\mathbb{R}}_{+}$ and $\beta:B\to{\mathbb{R}}_{+}$ satisfying properties (i),(ii) in that lemma; moreover, as is seen from its proof, for any number $\xi>0$, the functions $\alpha,\beta$ can be chosen so as to satisfy the equality $\sum(\alpha(e)\colon e\in A)=\xi$. Take $\xi:=\varepsilon$, define $\phi_{f}:=\alpha(e)$ for $f\in F_{K}$ and $e\in D_{f}$, and define $\psi_{w}:=\beta(e)$ for $w\in W_{K}$ and $e\in H_{w}$. These $\phi,\psi$ extended by zero outside $K$ satisfy (4.5)–(4.8), as required. $\Box$
Lemma 4.4 has a sharper version. To state it, consider the directed active graph $\overrightarrow{\Gamma}=(V_{\Gamma},\overrightarrow{E}_{\Gamma})$ (see Definition 4) and its strong components. (Recall that a strong component of a directed graph is a maximal subgraph $L=(V_{L},E_{L})$ of this graph such that $|V_{L}|>1$ and for any ordered pair $(u,v)$ in $V_{L}$, there is a directed path from $u$ to $v$. In other words, $L$ is a maximal connected subgraph representable as the union of directed cycles.) A weak component $K$ in $\overrightarrow{\Gamma}$ may contain several strong components; we denote the set of these by ${\cal L}(K)$. There is a natural partial order on ${\cal L}(K)$ in which $L\in{\cal L}(K)$ is regarded as preceding $L^{\prime}\in{\cal L}(K)$ if $K$ has a directed path going from $L$ to $L^{\prime}$. We denote by ${\cal L}^{\rm max}(K)$ the set of maximal components in ${\cal L}(K)$ (i.e., those having no succeeding components). Note that ((4.3)) implies that if $v$ is a vertex of $L\in{\cal L}^{\rm max}(K)$, then any edge of $K$ leaving $v$ belongs to $L$ as well.
We associate with $(\phi,\psi)\in{\mathbb{R}}_{+}^{F^{+}}\times{\mathbb{R}}_{+}^{W^{+}}$ the function $\overrightarrow{\Delta}=\overrightarrow{\Delta}^{\phi,\psi}$ on the set of active directed edges $\overrightarrow{E}_{\Gamma}$, taking the value $\phi_{f}$ on the edges $(f,w)\in D_{f}$, and $\psi_{w}$ on the edges $(w,f)\in H_{w}$ (cf. $\Delta^{\phi,\psi}$ defined in ((4.10))). This $\overrightarrow{\Delta}$ represents a circulation in $\overrightarrow{\Gamma}$, because the equalities in (4.5) are equivalent to the conservation (or balance) conditions $\overrightarrow{\Delta}(\delta^{\rm out}(f))=\overrightarrow{\Delta}(\delta^{\rm in}(f))$ for $f\in F^{+}$, and similarly for (4.6) and $W^{+}$. Here $\delta^{\rm in}(v)$ ($\delta^{\rm out}(v)$) denotes the set of edges in $\overrightarrow{\Gamma}$ entering (resp. leaving) a vertex $v$. The following fact is of use:
@@
    for $(\phi,\psi)\in{\cal P}$, if vertices $u,v$ are connected in $\overrightarrow{\Gamma}$ by a directed path from $u$ to $v$, and if $(\phi,\psi)_{u}>0$, then $(\phi,\psi)_{v}>0$ as well.
(It suffices to see this for a single edge $e=(u,v)$ in $\overrightarrow{\Gamma}$. This edge belongs to the head at $u$ (either $H_{u}$ or $D_{u}$, in the directed version), and if $(\phi,\psi)_{u}>0$, then $\overrightarrow{\Delta}^{\phi,\psi}(e)>0$, implying $(\phi,\psi)_{v}>0$, by the balance condition at $v$.)
@@
Let $K$ be a weak component of $\overrightarrow{\Gamma}$. Then for any maximal component $L\in{\cal L}^{\rm max}(K)$, there exists a solution $(\phi,\psi)\in{\mathbb{R}}_{+}^{F^{+}}\times{\mathbb{R}}_{+}^{W^{+}}$ to system (4.5)–(4.8) whose values outside $L$ are zeros, and this solution is unique.
Proof  Consider $L=(V_{L},E_{L})\in{\cal L}^{\rm max}(K)$. As is mentioned above, for each vertex $v\in V_{L}$, the set $\delta^{\rm out}(v)$ of edges of $\overrightarrow{\Gamma}$ leaving $v$ is contained in $L$. So we can apply Lemma 4.3 to the sets $A:=\cup(D_{f}\colon f\in F_{L})$ and $B:=\cup(H_{w}\colon w\in W_{L})$. This gives rise to a solution $(\phi,\psi)\in{\cal P}$ to (4.5)–(4.8) having the support $\{v\colon(\phi,\psi)_{v}\neq 0\}$ within $L$.
The see the uniqueness of $(\phi,\psi)$, we first observe that its support is exactly $V_{L}$. (This follows for the strong connectedness of $L$ and property ((4.14)).)
Now suppose, for a contradiction, that there is another solution $(\phi^{\prime},\psi^{\prime})\in{\cal P}$ having the support within $L$. Then one can choose numbers $\lambda,\mu>0$ such that the function $(\phi^{\prime\prime},\psi^{\prime\prime}):=(\lambda\phi,\lambda\psi)-(\mu\phi^{\prime},\mu\psi^{\prime})$ is nonnegative, satisfies (4.5)–(4.8), and takes zero value on some $f\in F_{L}$ or $w\in W_{L}$. But this is impossible by the observation above. $\Box$
@@
For each weak component $K$ of $\overrightarrow{\Gamma}$, any function $(\phi,\psi)\in{\mathbb{R}}_{+}^{F^{+}}\times{\mathbb{R}}_{+}^{W^{+}}$ satisfying (4.5)–(4.7) takes zero value on all vertices that belong to none of the maximal strong components in $K$.
Proof Let $Y$ be the set of vertices occurring in components in ${\cal L}^{\rm max}(K)$. Consider a function $(\phi,\psi)$ satisfying (4.5)–(4.7), and suppose, for a contradiction, that the set $Z$ of vertices $v\in V_{K}-Y$ with nonzero values $(\phi,\psi)_{v}$ is nonempty. Since the strong components in ${\cal L}^{\rm max}(K)$ are maximal, the set $\delta^{\rm out}(Y)$ of edges of $K$ going from $Y$ to $V_{K}-Y$ is empty.
On the other hand, for any vertex $u$ in $V_{K}-Y$, in particular, for $u\in Z$, there is a directed path $P$ going from $u$ to a maximal strong component in $K$. Then, by ((4.14)), $(\phi,\psi)_{u}>0$ implies $(\phi,\psi)_{v}>0$ for all vertices $v$ on this path; therefore, the edge $e$ of $P$ leaving $V_{K}-Y$ and entering $Y$ satisfies $\overrightarrow{\Delta}^{\phi,\psi}(e)>0$.
@@
where $\delta^{\rm in}(Y)$ is the set of edges of $K$ entering $Y$. But the difference in the left hand side must be zero (as this is the sum, over the vertices $v$ of $Y$, of “divergences” $\overrightarrow{\Delta}(\delta^{\rm in}(v))-\overrightarrow{\Delta}(\delta^{\rm out}(v))$, which are zeros). A contradiction. $\Box$
Let ${\cal L}^{\rm rot}={\cal L}^{\rm rot}(x)$ denote the set of maximal strong components $L$ occurring in the sets ${\cal L}^{\rm max}(K)$ over all weak components $K$ of the active graph $\overrightarrow{\Gamma}(x)$. By Lemma 4.5, for each component $L\in{\cal L}^{\rm rot}$, there is a unique solution to (4.5)–(4.8) having the support in $L$ (which is, moreover, exactly $V_{L}$); we denote this solution as $(\phi_{L}^{\varepsilon},\psi_{L}^{\varepsilon})=(\phi_{L}^{\varepsilon}(x),\psi_{L}^{\varepsilon}(x))$. Then Lemmas 4.5 and 4.6 imply
@@
The vertices of the polytope ${\cal P}^{\varepsilon}(x)$ (defined in (4.9)) are exactly the vectors $(\phi_{L}^{\varepsilon},\psi_{L}^{\varepsilon})$ for $L\in{\cal L}^{\rm rot}$. The supports of these vectors are pairwise disjoint.
@@
Definition 5. By a ${\mathbb{Z}}$-circulation we mean an integer-valued function $\rho:E\to{\mathbb{Z}}$ such that $\rho\neq 0$ and for each vertex $v\in V$, the sum $\rho(E_{v})$ ($=\sum(\rho(e)\colon e\in E_{v})$) is zero. Such a $\rho$ is called aligned if $\rho(fw),\rho(fw^{\prime})>0$ implies $\rho(fw)=\rho(fw^{\prime})$, and, symmetrically, $\rho(fw),\rho(f^{\prime}w)<0$ implies $\rho(fw)=\rho(f^{\prime}w)$. We say that $\rho$ is normalized if the g.c.d. of the values $\rho(e)$, $e\in E$, is one. We refer to a normalized aligned ${\mathbb{Z}}$-circulation $\rho$ for which the subgraph of $G$ induced by the support $\{e\in E\colon\rho(e)\neq 0\}$ of $\rho$ is connected as a pre-rotation in $G$.
Definition 6. Consider a maximal strong component $L\in{\cal L}^{\rm rot}$ and the corresponding vector $(\phi_{L}^{\varepsilon},\psi_{L}^{\varepsilon})\in{\cal P}^{\varepsilon}(x)$ with the support $V_{L}$ (existing by Lemma 4.5). Then there is a normalized aligned $Z$-circulation (viz. pre-rotation) proportional to the function $\Delta=\Delta^{\phi_{L}^{\varepsilon},\psi_{L}^{\varepsilon}}$ on $E$ defined as in ((4.10)); we denote this circulation by $\rho_{L}$ and call it the rotation for $x$ associated with $L$. Its analog for $\overrightarrow{\Gamma}$ is denoted as $\overrightarrow{\rho}_{L}$ (which for $\overrightarrow{e}\in\overrightarrow{E}_{\Gamma}$, takes the value $|\Delta(e)|$). We say that a weight $\lambda\in{\mathbb{R}}_{+}$ is admissible for $L$, or for $\rho_{L}$, under the assignment $x$ if the following holds:
@@
    (a) the function $x^{\prime}:=x+\lambda\rho_{L}$ is nonnegative and does not exceed the upper bound $b$, and (b) for each $w\in W^{+}$, either $\pi^{\rm c}_{w}(x^{\prime})=\pi^{\rm c}_{w}(x)$ and $H_{w}(x^{\prime})\supseteq H_{w}(x)$, or $x^{\prime}(e)=0$ for all edges $e$ in $\pi^{\rm c}_{w}(x)$ (in the latter case, $x(e)=\lambda$ is valid, and the critical tie at $w$ becomes better: $\pi^{\rm c}_{w}(x^{\prime})>\pi^{\rm c}_{w}(x)$).
The maximal admissible weight $\lambda$ for $L$ is denoted by $\tau_{L}=\tau_{L}(x)$ and called the maximal rotational weight for $L$, or for $\rho_{L}$.
Remark 1. It is well known that all rotations arising in stable allocation problems are $0,\pm 1$ functions whose supports are formed by simple cycles. In contrast, in SMP, rotations may take large values (even exponentially large in $|E|$). An instance of a “big” rotation in a “small” graph will be demonstrated in Example 3 in Sect. 5.
Note that Lemma 4.5 provides the existence of an admissible weight $\lambda>0$ for each $L\in{\cal L}^{\rm rot}$. Clearly if $\lambda$ is admissible, then any $\lambda^{\prime}$ in the interval $[0,\lambda]$ is admissible as well. Moreover, as a refinement of ((4.15)) in case $\lambda<\tau_{L}$, one can see that
@@
    if $\lambda<\tau_{L}$, then the transformation of $x$ into $x^{\prime}:=x+\lambda\rho_{L}$ preserves the active graph $\Gamma$ and: for each $w\in V^{+}$, we have $\pi^{\rm c}_{w}(x^{\prime})=\pi^{\rm c}_{w}(x)$ and $H_{w}(x^{\prime})=H_{w}(x)$, and for each $f\in F^{+}$, we have $H_{f}(x^{\prime})=D_{f}(x)$.
This implies that in case $\lambda<\tau_{L}$, the rotation $\rho_{L}$ remains applicable to the new $x^{\prime}$, but now the maximal weight for $\rho_{L}$ reduces to $\tau_{L}(x)-\lambda$ ($=\tau_{L}(x^{\prime})$).
@@
For any $0<\lambda\leq\tau_{L}$, the function $x^{\prime}:=x+\lambda\Delta_{L}$ is a stable assignment, and $x^{\prime}\prec x$.
Proof  We have $0\leq x^{\prime}\leq b$ (by ((4.15))), and $|x^{\prime}_{v}|=|x_{v}|$ for all $v\in V$ (by the definition of $\rho_{L}$). So $x^{\prime}$ is a rational assignment with the same sets $F^{=}$ and $W^{=}$.
To show the stability of $x^{\prime}$, consider an edge $e=fw\in E$ with $x^{\prime}(e)<b(e)$. We have to show that $e$ is not blocking for $x^{\prime}$, i.e., there is $v\in\{f,w\}$ such that the edge $e$ either belongs to the head $H_{v}(x^{\prime})$ or occurs in a tie in $\Pi^{\rm end}_{v}(x^{\prime})$; then $e$ is not interesting for $v$ under $x^{\prime}$, and we are done. (Here it suffices to consider the case when both $f,w$ are fully filled; for otherwise Cleaning procedure ensures that $x^{\prime}_{f}=x_{f}$ and $x^{\prime}_{w}=x_{w}$.)
To show this, we first suppose that $x^{\prime}(e)\neq x(e)$. The case $x^{\prime}(e)<x(e)$ is possible only if $w\in V_{L}$ and $e\in H_{w}(x)$. Then (cf. ((4.15))(b)) either $e\in H_{w}(x^{\prime})$ or $e\in\pi^{i}<\pi^{\rm c}_{w}(x^{\prime})$. And if $x^{\prime}(e)>x(e)$, then $f\in V_{L}$ and $e\in D_{f}(x)$, implying $e\in H_{f}(x^{\prime})$. Hence in both cases, $e$ is not blocking for $x^{\prime}$.
Now let $x^{\prime}(e)=x(e)<b(e)$. Then $e\notin E_{L}$. Since $e$ is not blocking for $x$ (as $x$ is stable), we are in at least one of the following cases: (i) $e\in H_{w}$; (ii) $e\in E^{\rm end}_{w}(x)$; (iii) $e\in H_{f}(x)$; (iv) $e\in E^{\rm end}_{f}(x)$ (where $w\in W^{=}$ in cases (i),(ii), and $f\in F^{=}$ in cases (iii),(iv)).

### v2

@@ 4 Rotations
(i) for each $v\in J$,  $\sum(\alpha(e)\colon e\in A_{v})=\sum(\beta(e)\colon e\in B_{v})$;
@@
Definition 4. Refining the term “active” mentioned above, we call an edge $e=fw$ active for $F$ if $f\in F^{+}$ and $e\in D_{f}$, and active for $W$ if $w\in W^{+}$ and $e\in H_{w}$. The subgraph of $G$ induced by the set of active edges is called the active graph and denoted as $\Gamma=\Gamma(x)=(V_{\Gamma},E_{\Gamma})$. The directed counterpart of the active graph, denoted as $\overrightarrow{\Gamma}=\overrightarrow{\Gamma}(x)=(V_{\Gamma},\overrightarrow{E}_{\Gamma})$, is obtained by replacing each edge $e=fw\in E_{\Gamma}$ by the directed edge $\overrightarrow{e}=(f,w)$ if $e\in D_{f}$, and by the directed edge $\overrightarrow{e}=(w,f)$ if $e\in H_{w}$. A (weak) component of $\Gamma$ (or $\overrightarrow{\Gamma}$) is a maximal subgraph $K$ in it such that any two vertices of $K$ are connected by a (not necessarily directed) path.
@@
Proof  Take as $A$ the set $\cup(D_{f}\colon f\in F_{K})$ and take as $B$ the set $\cup(H_{w}\colon w\in W_{K})$, where $F_{K}:=F\cap V_{K}$ and $W_{K}:=W\cap V_{K}$. Property ((4.5)) implies that $A,B$ obey condition ((4.13)). Then, by Lemma 4.5, there exist $\alpha:A\to{\mathbb{R}}_{+}$ and $\beta:B\to{\mathbb{R}}_{+}$ satisfying properties (i),(ii) in that lemma; moreover, as is seen from its proof, for any number $\xi>0$, the functions $\alpha,\beta$ can be chosen so as to satisfy the equality $\sum(\alpha(e)\colon e\in A)=\xi$. Take $\xi:=\varepsilon/2$, define $\phi_{f}:=\alpha(e)$ for $f\in F_{K}$ and $e\in D_{f}$, and define $\psi_{w}:=\beta(e)$ for $w\in W_{K}$ and $e\in H_{w}$. These $\phi,\psi$ extended by zero outside $K$ satisfy (4.7)–(4.10), as required. $\Box$
Lemma 4.6 has a sharper version. To state it, consider the directed active graph $\overrightarrow{\Gamma}=(V_{\Gamma},\overrightarrow{E}_{\Gamma})$ (see Definition 4), and let ${\cal L}={\cal L}(x)$ denote the set of strong components in $\overrightarrow{\Gamma}$. (Recall that a strong component of a directed graph is a maximal (by inclusion) subgraph $L=(V_{L},E_{L})$ in it such that $|V_{L}|>1$ and for any ordered pair $(u,v)$ in $V_{L}$, there is a directed path from $u$ to $v$. In other words, $L$ is a maximal connected subgraph representable as the union of directed cycles.) We write $F_{L}$ for $V_{L}\cap F$ and $W_{L}$ for $V_{L}\cap W$. There is a natural partial order on ${\cal L}$ in which $L\in{\cal L}$ is regarded as preceding $L^{\prime}\in{\cal L}$ if $\overrightarrow{\Gamma}$ has a directed path going from $L$ to $L^{\prime}$. We denote by ${\cal L}^{\rm rot}={\cal L}^{\rm rot}(x)$ the set of maximal components in ${\cal L}$ under this order (i.e., those having no succeeding components). Note that ((4.5)) implies that if $v$ is a vertex of $L\in{\cal L}^{\rm rot}$, then any edge of $\overrightarrow{\Gamma}$ leaving $v$ belongs to $L$ as well.
We associate with $(\phi,\psi)\in{\mathbb{R}}_{+}^{F^{+}}\times{\mathbb{R}}_{+}^{W^{+}}$ the function $\overrightarrow{\Delta}=\overrightarrow{\Delta}^{\phi,\psi}$ on the set of active directed edges $\overrightarrow{E}_{\Gamma}$, taking the value $\phi_{f}$ on the edges $(f,w)\in D_{f}$, and $\psi_{w}$ on the edges $(w,f)\in H_{w}$ (Compare with $\Delta^{\phi,\psi}$ defined in ((4.12))). When $(\phi,\psi)\in{\cal P}^{\varepsilon}$, this $\overrightarrow{\Delta}$ represents a circulation in $\overrightarrow{\Gamma}$, because the equalities in (4.7) are equivalent to the conservation (or balance) conditions $\overrightarrow{\Delta}(\delta^{\rm out}(f))=\overrightarrow{\Delta}(\delta^{\rm in}(f))$ for $f\in F^{+}$, and similarly for (4.8) and $W^{+}$. Here $\delta^{\rm in}(v)$ ($\delta^{\rm out}(v)$) denotes the set of edges in $\overrightarrow{\Gamma}$ entering (resp. leaving) a vertex $v$. The following fact is of use:
@@
    for $(\phi,\psi)\in{\cal P}^{\varepsilon}$, if vertices $u,v$ are connected in $\overrightarrow{\Gamma}$ by a directed path from $u$ to $v$, and if $(\phi,\psi)_{u}>0$, then $(\phi,\psi)_{v}>0$ as well.
(It suffices to see this for a single edge $e=(u,v)$ in $\overrightarrow{\Gamma}$. This edge belongs to the active head at $u$ (the directed version of either $H_{u}$ or $D_{u}$), and if $(\phi,\psi)_{u}>0$, then $\overrightarrow{\Delta}^{\phi,\psi}(e)>0$, implying $(\phi,\psi)_{v}>0$ by the balance condition at $v$.)
@@
For any maximal strong component $L\in{\cal L}^{\rm rot}$, there exists a solution $(\phi,\psi)\in{\mathbb{R}}_{+}^{F^{+}}\times{\mathbb{R}}_{+}^{W^{+}}$ to system (4.7)–(4.10) whose values outside $L$ are zeros, and this solution is unique.
Proof  Consider $L=(V_{L},E_{L})\in{\cal L}^{\rm rot}$. As is mentioned above, for each vertex $v\in V_{L}$, the set $\delta^{\rm out}(v)$ of edges of $\overrightarrow{\Gamma}$ leaving $v$ is contained in $L$. So we can apply Lemma 4.5 to the sets $A:=\cup(D_{f}\colon f\in F_{L})$ and $B:=\cup(H_{w}\colon w\in W_{L})$. This gives rise to a solution $(\phi,\psi)\in{\cal P}^{\varepsilon}$ to (4.7)–(4.10) having the support $\{v\colon(\phi,\psi)_{v}\neq 0\}$ within $L$.
To see the uniqueness of $(\phi,\psi)$, we first observe that its support is exactly $V_{L}$. (This follows for the strong connectedness of $L$ and property ((4.15)).)
Now suppose, for a contradiction, that there is another solution $(\phi^{\prime},\psi^{\prime})\in{\cal P}^{\varepsilon}$ having the support within $L$. Then one can choose $\lambda,\mu>0$ such that the function $(\phi^{\prime\prime},\psi^{\prime\prime}):=(\lambda\phi,\lambda\psi)-(\mu\phi^{\prime},\mu\psi^{\prime})$ is nonnegative, satisfies (4.7)–(4.10), and takes zero value on some vertex $f\in F_{L}$ or $w\in W_{L}$. But this is impossible by the observation above. $\Box$
@@
Any function $(\phi,\psi)\in{\mathbb{R}}_{+}^{F^{+}}\times{\mathbb{R}}_{+}^{W^{+}}$ satisfying (4.7),(4.8),(4.10) takes zero value on all vertices not contained in maximal strong components of $\overrightarrow{\Gamma}$.
Proof  Let $Y$ be the set of vertices occurring in components in ${\cal L}^{\rm rot}$. Consider a function $(\phi,\psi)$ satisfying (4.7),(4.8),(4.10), and suppose, for a contradiction, that the set $Z$ of vertices $v\in V_{\Gamma}-Y$ with nonzero values $(\phi,\psi)_{v}$ is nonempty. Since ${\cal L}^{\rm rot}$ consists of maximal components, the set $\delta^{\rm out}(Y)$ of edges of $\overrightarrow{\Gamma}$ going from $Y$ to $V_{\Gamma}-Y$ is empty.
On the other hand, for any vertex $u$ in $V_{\Gamma}-Y$, in particular, for $u\in Z$, there is a directed path $P$ going from $u$ to a maximal strong component of $\overrightarrow{\Gamma}$. Then, by ((4.15)), $(\phi,\psi)_{u}>0$ implies $(\phi,\psi)_{v}>0$ for all vertices $v$ on this path; therefore, the edge $e$ of $P$ leaving $V_{\Gamma}-Y$ and entering $Y$ satisfies $\overrightarrow{\Delta}^{\phi,\psi}(e)>0$.
@@
where $\delta^{\rm in}(Y)$ is the set of edges of $\overrightarrow{\Gamma}$ entering $Y$. But the difference in the left hand side must be zero (as this is the sum, over the vertices $v$ of $Y$, of “divergences” $\overrightarrow{\Delta}(\delta^{\rm in}(v))-\overrightarrow{\Delta}(\delta^{\rm out}(v))$, which are zeros). A contradiction. $\Box$
By Lemma 4.7, for each component $L\in{\cal L}^{\rm rot}$, there is a unique solution to (4.7)–(4.10) having the support in $L$ (which is, moreover, exactly $V_{L}$); we denote this solution as $(\phi_{L}^{\varepsilon},\psi_{L}^{\varepsilon})=(\phi_{L}^{\varepsilon}(x),\psi_{L}^{\varepsilon}(x))$. Then Lemmas 4.7 and 4.8 imply
@@
The vertices of polytope ${\cal P}^{\varepsilon}(x)$ (defined in (4.11)) are exactly the vectors $(\phi_{L}^{\varepsilon},\psi_{L}^{\varepsilon})$ for $L\in{\cal L}^{\rm rot}(x)$. The supports of these vectors are pairwise disjoint.
@@
Definition 5. By a ${\mathbb{Z}}$-circulation we mean an integer-valued function $\rho:E\to{\mathbb{Z}}$ such that $\rho\neq 0$ and for each vertex $v\in V$, the sum $\rho(E_{v})$ ($=\sum(\rho(e)\colon e\in E_{v})$) is zero. Such a $\rho$ is called aligned if $\rho(fw),\rho(fw^{\prime})>0$ implies $\rho(fw)=\rho(fw^{\prime})$, and, symmetrically, $\rho(fw),\rho(f^{\prime}w)<0$ implies $\rho(fw)=\rho(f^{\prime}w)$. We say that $\rho$ is normalized if the g.c.d. of values $\rho(e)$, $e\in E$, is one. We refer to a normalized aligned ${\mathbb{Z}}$-circulation $\rho$ for which the subgraph of $G$ induced by the support $\{e\in E\colon\rho(e)\neq 0\}$ of $\rho$ is connected as a pre-rotation in $G$.
Definition 6. Consider a maximal strong component $L\in{\cal L}^{\rm rot}$ and the corresponding vector $(\phi_{L}^{\varepsilon},\psi_{L}^{\varepsilon})\in{\cal P}^{\varepsilon}(x)$ with the support $V_{L}$ (existing by Lemma 4.7). Then there is a normalized aligned $Z$-circulation (viz. pre-rotation) proportional to the function $\Delta=\Delta^{\phi_{L}^{\varepsilon},\psi_{L}^{\varepsilon}}$ on $E$ defined as in ((4.12)); we denote this circulation by $\rho_{L}$ and call it the rotation for $x$ associated with $L$. Its analog for $\overrightarrow{\Gamma}$ is denoted as $\overrightarrow{\rho}_{L}$ (which for each $\overrightarrow{e}\in\overrightarrow{E}_{\Gamma}$, takes the value $|\Delta(e)|$). We say that a weight $\lambda\in{\mathbb{R}}_{+}$ is admissible for $L$, or for $\rho_{L}$, under the assignment $x$ if the following holds:
@@
    (a) the function $x^{\prime}:=x+\lambda\rho_{L}$ is nonnegative and does not exceed the upper bound $b$, and (b) for each $w\in W^{+}$, either $\pi^{\rm c}_{w}(x^{\prime})=\pi^{\rm c}_{w}(x)$ and $H_{w}(x^{\prime})\supseteq H_{w}(x)$, or $x^{\prime}(e)=0$ for all edges $e$ in $\pi^{\rm c}_{w}(x)$ (in the latter case, $x(e)=\lambda$, and the critical tie at $w$ becomes better: $\pi^{\rm c}_{w}(x^{\prime})>\pi^{\rm c}_{w}(x)$).
The maximal admissible weight $\lambda$ for $L$ is denoted by $\tau_{L}(x)$ and called the maximal rotational weight for $L$, or for $\rho_{L}$, under $x$.
Remark 1. It is well known that all rotations arising in instances of stable allocation problem (SAP) are $0,\pm 1$ functions whose supports are formed by simple cycles. In contrast, in SMP, rotations may take large values (even exponentially large in $|E|$). A construction of “big” rotations will be given in Example 3 and Remark 3 in Sect. 5.
Lemma 4.7 provides the existence of an admissible weight $\lambda>0$ for each $L\in{\cal L}^{\rm rot}$. Clearly if $\lambda$ is admissible, then any $\lambda^{\prime}$ in the interval $[0,\lambda]$ is admissible as well. Moreover, as a refinement of ((4.16)) in case $\lambda<\tau_{L}$, one can see that
@@
    if $\lambda<\tau_{L}(x)$, then the transformation of $x$ into $x^{\prime}:=x+\lambda\rho_{L}$ preserves the active graph $\Gamma$ and: for each $w\in V^{+}$, we have $\pi^{\rm c}_{w}(x^{\prime})=\pi^{\rm c}_{w}(x)$ and $H_{w}(x^{\prime})=H_{w}(x)$, and for each $f\in F^{+}$, we have $H_{f}(x^{\prime})=D_{f}(x)$.
This implies that in case $\lambda<\tau_{L}(x)$, the rotation $\rho_{L}$ remains applicable to the new $x^{\prime}$, but now the maximal weight for $\rho_{L}$ reduces to $\tau_{L}(x)-\lambda$ ($=\tau_{L}(x^{\prime})$).
@@
For any $\lambda\leq\tau_{L}(x)$, the assignment $x^{\prime}:=x+\lambda\Delta_{L}$ is stable and $x^{\prime}\prec x$.
Proof  We have $0\leq x^{\prime}\leq b$ (by ((4.16))), and $|x^{\prime}_{v}|=|x_{v}|$ for all $v\in V$ (by the definition of $\rho_{L}$). So $x^{\prime}$ obeys both $b$ and $q$ and preserves the fully filled sets $F^{=}$ and $W^{=}$.
To show the stability of $x^{\prime}$, consider an edge $e=fw\in E$ with $x^{\prime}(e)<b(e)$. We have to show that $e$ is not blocking for $x^{\prime}$; equivalently, there is $v\in\{f,w\}$ such that the edge $e$ either belongs to the head $H_{v}(x^{\prime})$ or occurs in $E^{\rm end}_{v}(x^{\prime})$. Then $e$ is not interesting for $v$ under $x^{\prime}$, and we are done. (Here it suffices to consider the case when both $f,w$ are fully filled; for otherwise Cleaning procedure (C) ensures that $x^{\prime}_{f}=x_{f}$ and $x^{\prime}_{w}=x_{w}$.)
To show this, we first suppose that $x^{\prime}(e)\neq x(e)$. The case $x^{\prime}(e)<x(e)$ is possible only if $w\in V_{L}$ and $e\in H_{w}(x)$. Then (cf. ((4.16))(b)) either $e\in H_{w}(x^{\prime})$ or $e\in\pi^{\rm c}_{w}(x)<\pi^{\rm c}_{w}(x^{\prime})$ (whence $e\in E^{\rm end}_{w}(x^{\prime})$). And if $x^{\prime}(e)>x(e)$, then $f\in V_{L}$ and $e\in D_{f}(x)$, implying $e\in H_{f}(x^{\prime})$. Hence in both cases, $e$ is not blocking for $x^{\prime}$.
Now let $x^{\prime}(e)=x(e)<b(e)$. Then $e\notin E_{L}$. Since $x$ is stable, we are in at least one of the following cases: (i) $e\in H_{w}(x)$; (ii) $e\in E^{\rm end}_{w}(x)$; (iii) $e\in H_{f}(x)$; (iv) $e\in\pi^{i}\subseteq E^{\rm end}_{f}(x)$ (where $w\in W^{=}$ in cases (i),(ii), and $f\in F^{=}$ in cases (iii),(iv)).